OpenStudy (anonymous):

Solve 15x^2 − 15x − 30 = 0 by completing the square. a) What is the missing term required to form a square trinomial?

OpenStudy (anonymous):

Step 1: Factor left side of equation. 15(x+1)(x−2)=0 Step 2: Set factors equal to 0. x+1=0 or x−2=0 x=−1 or x=2 Answer: x=−1 or x=2